OpenStudy (anonymous):

Which of the following is an equation of a line with a slope of -2 that passes through the point (-4,3)? Help.

OpenStudy (anonymous):

Could you show us the following?

OpenStudy (anonymous):

I just type the question it ask me

OpenStudy (anonymous):

?? in your question, you ask pick one of the following. So because there is no possible answers for us to choose from we cannot answer your question.

OpenStudy (anonymous):

Here the answer choices A.y=-2x-5 B.y=-2x-4 C.y=-2x+3 D.y=-2x+11 that all it showing on my dame paper man im just trying to get some help before i take the ECA's tomorrow

OpenStudy (anonymous):

So if you ain't gone help me please leave

OpenStudy (anonymous):

Ok. So the easiest way to solve this problem is to simply "plug" in the answers. So the first number in ur () is ur x and the second ur y. So find the one that works. In this case the answer is d

OpenStudy (anonymous):

Ok how did you get that answer show me the work you did im a hands on learning guy i can't just go by what you tell me i need to see it so draw it out of something please

OpenStudy (anonymous):

whoops my bad wrong answer . correct answer is a

OpenStudy (anonymous):

Ill show u it in a sec

OpenStudy (anonymous):



OpenStudy (anonymous):

So how did you know that it was A

OpenStudy (anonymous):

I simply plugged it in in my head. As you progress in math, previous things become easier. In about a year or two it should be the same for u. Its like when you were learning adding and subtracting it took some time. but if you do it now its simple. Be safe and write it down for now tho espec for a test.

OpenStudy (anonymous):

Oh well i mean what about the others why didnt you pick them

OpenStudy (anonymous):

Oh.. i did. i plugged them all in.

OpenStudy (tamtoan):

jaydanzy, the best thing is to find the equation of a line with the given information then compare your answer with what they gave you. equation of a line will have the form ...y = mx + b ...where m is the slope. so your equation will be in the form of: y = -2x + b <--- you need to find b. since the line pass through point (-4,3), substitute this point into the equation, you have: 3 = -2(-4) + b = 8 + b ===> b = -5 now you have all the info. your final answer is : y = -2x - 5 ...

OpenStudy (anonymous):

I don't get it but thanks anyway